OpenStudy (anonymous):

Find the inverse of y = 7x - 3

OpenStudy (anonymous):

@Harsha19111999 @TheEdwardsFamily @IrishBoy123 @Godlovesme @kalee.schaefer @logan420 @phi

OpenStudy (anonymous):

what you think?

OpenStudy (anonymous):

Reverse?

OpenStudy (anonymous):

are there any options?

OpenStudy (anonymous):

@iGreen can you help me?

OpenStudy (anonymous):

@iGreen.

OpenStudy (amorfide):

if you want to find the inverse that means you want to make x the subject so y=7x-3 y+3=7x (y+3)/7=x replace x to be \[f ^{-1}(x)\] and replace y with x to get \[f^{-1}(x)=\frac{ x+3 }{ 7 }\]

OpenStudy (harsha19111999):

Inverse in what sense? It may be additive inverse or multiplicative inverse or function inverse. For example, the additive inverse of 2 is -2 and the multiplicative inverse of 2 is 1/2 and for functions, its different

OpenStudy (anonymous):

inverse f(x) =(x+3)/7